# Dissecting Mahalanobis: How Feature Geometry and Normalization Shape OOD Detection

### v1

@@ D.2 Detailed OOD performance
Table 4: False-positive rate (FPR, $\downarrow$) across models for NINCO dataset using different Mahalanobis variants: MD* uses the empirically optimal $\beta$, $\hat{\textbf{MD}}$ uses the regression-predicted $\hat{\beta}$, MD (standard) fixes $\beta=0$, and MD++ (Mahalanobis++) fixes $\beta=1$.
Model MD* $\hat{\text{MD}}$ MD MD++ RMD* $\hat{\text{RMD}}$ RMD RMD++
BEiTV2 FT In1k 0.446 0.463 0.506 0.470 0.414 0.418 0.457 0.443
BEiTV2 FT In21k 0.281 0.281 0.475 0.364 0.311 0.322 0.339 0.331
DINOV2 0.403 0.415 0.424 0.445 0.573 0.573 0.577 0.574
DeiT3 0.496 0.496 0.505 0.500 0.422 0.429 0.437 0.433
DeiT3 FB In22k In1k 0.455 0.464 0.480 0.455 0.445 0.455 0.457 0.445
DeiT3 In21k In1k 0.423 0.426 0.432 0.423 0.349 0.349 0.407 0.387
DeiT3-L In22k In1k 0.386 0.391 0.402 0.388 0.368 0.368 0.405 0.385
EVA02 0.535 0.630 0.691 0.630 0.499 0.515 0.515 0.523
EVA02 FT In1k 0.409 0.412 0.418 0.413 0.418 0.418 0.431 0.425
EVA02 FT In21k 0.538 0.546 0.638 0.563 0.385 0.393 0.393 0.400
EVA02 FT In21k In1k 0.406 0.406 0.424 0.407 0.387 0.388 0.425 0.406
EVA02-L FT In22k In1k 0.349 0.358 0.356 0.388 0.311 0.312 0.321 0.316
EVA02-S FT In22k In1k 0.494 0.499 0.526 0.504 0.506 0.506 0.517 0.513
MAE FT In1k 0.481 0.485 0.516 0.488 0.428 0.433 0.463 0.450
ViT 0.553 0.557 0.557 0.556 0.511 0.513 0.516 0.514
ViT CLIP In12k In1k 0.302 0.317 0.379 0.336 0.301 0.301 0.373 0.352
ViT CLIP In1k 0.436 0.439 0.476 0.455 0.409 0.411 0.446 0.433
ViT In21K 0.630 0.659 0.641 0.658 0.442 0.454 0.451 0.453
ViT In21K In1k 0.394 0.394 0.405 0.481 0.415 0.422 0.425 0.425
ViT-L CLIP In12k In1k 0.309 0.309 0.320 0.310 0.272 0.282 0.299 0.290
ViT-L In21K In1k 0.303 0.306 0.322 0.358 0.264 0.272 0.272 0.277
ViT-S In21K In1k 0.499 0.500 0.515 0.505 0.512 0.514 0.512 0.514
Appendix E Geometry of Eigvenvalues
@@
Appendix F Pearson correlations
@@
Appendix G Full model names

### v3

@@ F.2 Detailed OOD performance
Table 7: FPR on NINCO across model families for Mahalanobis variants (lower is better). MD* uses the empirically optimal $\beta$; $\hat{\textbf{MD}}$ uses the regression-predicted $\hat{\beta}$; MD (standard) fixes $\beta=0$; and MD++ (Mahalanobis++) fixes $\beta=1$.
Model Acc MSP MLS KNN VIM RS-MD MD++ MD RS-RMD RMD++ RMD
[…]
BEiTV2 FT In21k 85.1 40.8 28.7 44.3 33.0 36.4 36.4 47.5 33.1 33.1 33.9
DINOV2 83.0 48.3 37.2 54.7 35.2 42.4 44.5 42.4 57.7 57.4 57.7
[…]
ViT-S In21K In1k 75.8 60.3 54.1 60.1 52.0 53.1 50.5 51.5 51.3 51.4 51.2
[…]
EVA02 FT In21k 80.2 47.1 45.8 56.5 40.7 56.3 56.3 63.8 40.0 40.0 39.3
[…]
DeiT3 FB In22k In1k 83.8 63.8 82.0 47.7 57.8 47.8 45.5 48.0 44.5 44.5 45.7
Avg 55.8 63.5 49.3 51.5 44.5 45.0 46.5 41.8 42.1 42.8
Appendix G Geometry of Eigvenvalues
@@
Appendix H Pearson correlations
@@
Appendix I Full model names